OpenStudy (anonymous):

0.5 x(times) (2n+6) =4n-12

OpenStudy (anonymous):

I dont understand what the question is asking?

OpenStudy (anonymous):

\[0.5\times(2n+6) = 5n - 12- n\]

OpenStudy (anonymous):

Do you know how to simplify the left side?

OpenStudy (lgbasallote):

0.5 means 1/2 so \[\frac{1}{2} (2n + 6) = 4n - 12\] distribute... \[ n + 3 = 4n - 12\] i guess you can solve now?

OpenStudy (anonymous):

thank you :>